OpenStudy (anonymous):

How do i graph: x – 3y = –12 2x – y = 1

OpenStudy (anonymous):

sorry. its x-3y= 12 2x-y=1

OpenStudy (anonymous):

you looking to graph both and find where they intersect?

OpenStudy (earthcitizen):

\[y = (x+12)/3,y = 2 x-1\]

OpenStudy (anonymous):

if you are going to do it by hand, probably the best think to do if fine the intercepts

OpenStudy (anonymous):

so for \[x-3y=12\] find the x intercept by setting y = 0 get \[x=12\] so you know (12,0) is on the graph. for the y - intercept set x = 0 and get \[-3y=12\implies y=-4\] so you know (0,-4) is on the graph. plot those two points, connect the dots

OpenStudy (anonymous):

repeat the process for \[2x-y=1\] or else rewrite as earthcitizen wrote as \[y=2x-1\] and then replace x by some well chosen numbers and find a couple points

OpenStudy (anonymous):

did this help?

OpenStudy (anonymous):

yes it di. thankS!